### v2

@@ 3.7 Forecasting in manifold space and baseline comparison
    Tangent PCA: Project $v_{t}$ on top-$d$ principal axes to obtain coefficients $c_{t}\in\mathbb{R}^{d}$.Jolliffe2002
@@
We explicitly compare against a geometry-agnostic baseline that applies the same predictors directly in the input space: fit VAR to the raw $\mathbb{R}^{3}$ series $\{X_{t}\}$ (or its first differences) without manifold embeddings or tangent/PCA steps, and produce $\hat{X}^{\text{native}}_{t+1:t+h}$. All training windows, horizons, and hyperparameters are matched to ensure a fair comparison. LopezDePrado2018AFML; Jansen2023; Jolliffe2002
@@ 3.8 Euclidean Null Control: Correlated Brownian Motions
Thus $\{W_{t}\}$ is a multivariate Brownian motion in flat $\mathbb{R}^{n}$ with constant diffusion $\Sigma_{\rho}$ (no curvature, no manifold constraints).RevuzYor1999
@@
revealing one dominant “market” factor and n1n{-}1 equal idiosyncratic directions. Projecting trajectories onto the top 2–3 PCs is a linear rotation/scale; it does not induce curvature.Jolliffe2002 Hence, if our method were merely “learning PCA’s geometry,” it would incorrectly report non-flat geometry here. The null control checks that it does not.
@@ 3.9 Real-Finance Data Pipeline: Expanding PCA, Eigenportfolios, and Forecasting
Let $R_{t}\in\mathbb{R}^{N}$ denote the cross-section of asset returns at time $t$. For $t\geq t_{0}$, we standardize the history $\{R_{1},\dots,R_{t}\}$ and compute PCA loadings $\{u_{1,t},u_{2,t},u_{3,t}\}$ and eigenvalues $\{\lambda_{1,t},\lambda_{2,t},\lambda_{3,t}\}$ on an expanding window (Jolliffe2002; Avellaneda2022; LopezDePrado2020MLAM. To avoid look-ahead, eigenportfolio $i$ at time $t$ uses the lagged loadings:
@@
Translating forecasts to trading signals and PnL.Maillard2010; Roncalli2013
@@ 3.10 Eigenvalue-Weighted Sleeves from 3D PCA Space (Expanding SVD)
where the coordinates are the out-of-sample cumulative eigenportfolio sleeves (PC1, PC2, PC3). Then, within this 3D space, we compute an expanding-window SVD/PCA to obtain time-varying variance levels and use those 3D-space eigenvalues to weight the sleeves,Jolliffe2002.
@@
All evaluation metrics (MAE/RMSE/Sign, Sharpe, cumulative PnL) are computed identically across (i) and (ii) to ensure a fair comparison; results are reported in Section 4.SimonianLopezdePradoFabozzi2024
@@ 4.3 Topological validation via persistent homology
We provide a topological check for toroidal behavior in the real-data embedding by estimating the homology of a delay-embedded attractor (Takens’ embedding, Takens1981) built from the 3D PCA path (and, where informative, from individual sleeves). The exhibit comprises: (i) a Takens attractor plot and (ii) persistent homology summaries (diagrams/barcodes) up to $H_{2}$. The torus test is heuristic but informative: a $T^{2}$-like attractor typically shows one connected component ($\beta_{0}\!=\!1$), two long-lived 1-cycles ($\beta_{1}\!\approx\!2$), and, in dense surface samples, a weak 2-class ($\beta_{2}\!\approx\!1$). Trajectory samples often recover the two $H_{1}$ classes cleanly, with $H_{2}$ less stable.
@@ 4.4 Trading Translation and PnL Construction
where $p_{i,t+1}$ is the out-of-sample eigenportfolio return and $\lambda_{i}(t)$ are the expanding-window 3D-space eigenvalues.Bailey2014DSR; Bailey2014PBO
@@ 4.5 Simulated Regimes: Cross–Manifold Performance
Table 1: Cross–manifold performance on simulated data,Hsu2002; doCarmoRiemannian. Best MAE/RMSE/MAPE and Sharpe per data–manifold block highlighted.
@@ 4.6 Validation on Correlated Brownian Motions – CBMs (Euclidean Null)
4.7 Financial Data: Coordinate Sharpe by Geometry,Maillard2010; Roncalli2013
@@
To put these results in context, we compare the geometry-informed strategies against conventional benchmark portfolios: the equally weighted long only (LO) portfolio of all assets, and a risk parity portfolio (RP) that balances contributions to volatility. Over the same 2005 to 2025 period, the equally weighted portfolio achieved a Sharpe ratio of roughly 0.39 and the RP portfolio about 0.44. The standout is the toroidal $z$ coordinate (Sharpe $0.721$), consistent with a latent cyclic component amplified by a $T^{2}$ embedding. In other words, a trading strategy informed by the toroidal geometry signals would have outperformed a passive diversified portfolio, highlighting the practical value of the geometric approach. Notably, even the spherical model’s Sharpe ($0.27$) is on par with the benchmarks, while the Euclidean forecast model’s negative Sharpe is clearly inferior. These findings reinforce that embedding financial time series in an appropriate curved space can extract predictive signals that traditional methods overlook. (Summary table of best performing strategies vs benchmarks in Table 3,Avellaneda2022)
@@ 5.1 IS–LM foundations for (multi-)torus dynamics: a mathematical sketch
1. IS–LM(+Phillips) linear core and oscillatory modes,Hicks1937
@@
The $(y,r)$-subsystem has complex eigenvalues (oscillatory adjustment,Benhabib1979; Barnett2008) iff
@@
Note on “Hopf”.Kuznetsov2013
@@ 3. From one torus to several: a multi-block macro
Writing $z^{(s)}=\rho_{s}e^{i\theta^{(s)}}$ and projecting on the limit cycles ($\rho_{s}\approx\rho_{s}^{\star}$), we obtain a phase network,Ikeda2012,
@@ 5.2 Future research agenda (testable directions)
Combine GIM with mixed-frequency nowcasting: update $(\theta_{t},\phi_{t})$ and the tangent-space forecasts as high-frequency data arrive, and reweight sleeves by expanding-SVD eigenvalues and curvature gates. This evaluates whether real-time macro signals enhance geometry-aware timing.NoguerAlonso2021FinEAS
@@ Appendix C Scenario catalog
Appendix D Coordinate transforms,doCarmoCurves; doCarmoRiemannian

### v3

@@ 4.7 Forecasting in manifold space and baseline comparison
    Tangent PCA: Project $v_{t}$ on top-$d$ principal axes to obtain coefficients $c_{t}\in\mathbb{R}^{d}$.Jolliffe (2002)
@@
We explicitly compare against a geometry-agnostic baseline that applies the same predictors directly in the input space: fit VAR to the raw $\mathbb{R}^{3}$ series $\{X_{t}\}$ (or its first differences) without manifold embeddings or tangent/PCA steps, and produce $\hat{X}^{\text{native}}_{t+1:t+h}$. All training windows, horizons, and hyperparameters are matched to ensure a fair comparison. López de Prado (2018); Jansen (2023); Jolliffe (2002)
@@ 4.8 Euclidean Null Control: Correlated Brownian Motions
Thus $\{W_{t}\}$ is a multivariate Brownian motion in flat $\mathbb{R}^{n}$ with constant diffusion $\Sigma_{\rho}$ (no curvature, no manifold constraints).Revuz and Yor (1999)
@@
revealing one dominant “market” factor and n1n{-}1 equal idiosyncratic directions. Projecting trajectories onto the top 2–3 PCs is a linear rotation/scale; it does not induce curvature.Jolliffe (2002) Hence, if our method were merely “learning PCA’s geometry,” it would incorrectly report non-flat geometry here. The null control checks that it does not.
@@ 4.9 Real-Finance Data Pipeline: Expanding PCA, Eigenportfolios, and Forecasting
Let $R_{t}\in\mathbb{R}^{N}$ denote the cross-section of asset returns at time $t$. For $t\geq t_{0}$, we standardize the history $\{R_{1},\dots,R_{t}\}$ and compute PCA loadings $\{u_{1,t},u_{2,t},u_{3,t}\}$ and eigenvalues $\{\lambda_{1,t},\lambda_{2,t},\lambda_{3,t}\}$ on an expanding window (Jolliffe (2002); Avellaneda et al. (2022); López de Prado (2020). To avoid look-ahead, eigenportfolio $i$ at time $t$ uses the lagged loadings:
@@
Translating forecasts to trading signals and PnL.Maillard et al. (2010); Roncalli (2013)
@@ 4.10 Eigenvalue-Weighted Sleeves from 3D PCA Space (Expanding SVD)
where the coordinates are the out-of-sample cumulative eigenportfolio sleeves (PC1, PC2, PC3). Then, within this 3D space, we compute an expanding-window SVD/PCA to obtain time-varying variance levels and use those 3D-space eigenvalues to weight the sleeves,Jolliffe (2002).
@@
All evaluation metrics (MAE/RMSE/Sign, Sharpe, cumulative PnL) are computed identically across (i) and (ii) to ensure a fair comparison; results are reported in Section 5.Simonian et al. (2024)
@@ 5.3 Topological validation via persistent homology
We provide a topological check for toroidal behavior in the real-data embedding by estimating the homology of a delay-embedded attractor (Takens’ embedding, Takens (1981)) built from the 3D PCA path (and, where informative, from individual sleeves). The exhibit comprises: (i) a Takens attractor plot and (ii) persistent homology summaries (diagrams/barcodes) up to $H_{2}$. The torus test is heuristic but informative: a $T^{2}$-like attractor typically shows one connected component ($\beta_{0}\!=\!1$), two long-lived 1-cycles ($\beta_{1}\!\approx\!2$), and, in dense surface samples, a weak 2-class ($\beta_{2}\!\approx\!1$). Trajectory samples often recover the two $H_{1}$ classes cleanly, with $H_{2}$ less stable.
@@ 5.4 Trading Translation and PnL Construction
where $p_{i,t+1}$ is the out-of-sample eigenportfolio return and $\lambda_{i}(t)$ are the expanding-window 3D-space eigenvalues.Bailey et al. (2014a, b)
@@ 5.5 Simulated Regimes: Cross–Manifold Performance
Table 1: Cross–manifold performance on simulated data,Hsu (2002a); do Carmo (1992). Best MAE/RMSE/MAPE and Sharpe per data–manifold block highlighted.
@@ 5.6 Validation on Correlated Brownian Motions – CBMs (Euclidean Null)
5.7 Financial Data: Coordinate Sharpe by Geometry,Maillard et al. (2010); Roncalli (2013)
@@
To put these results in context, we compare the geometry-informed strategies against conventional benchmark portfolios: the equally weighted long only (LO) portfolio of all assets, and a risk parity portfolio (RP) that balances contributions to volatility. Over the same 2005 to 2025 period, the equally weighted portfolio achieved a Sharpe ratio of roughly 0.39 and the RP portfolio about 0.44. The standout is the toroidal $z$ coordinate (Sharpe $0.721$), consistent with a latent cyclic component amplified by a $T^{2}$ embedding. In other words, a trading strategy informed by the toroidal geometry signals would have outperformed a passive diversified portfolio, highlighting the practical value of the geometric approach. Notably, even the spherical model’s Sharpe ($0.27$) is on par with the benchmarks, while the Euclidean forecast model’s negative Sharpe is clearly inferior. These findings reinforce that embedding financial time series in an appropriate curved space can extract predictive signals that traditional methods overlook. (Summary table of best performing strategies vs benchmarks in Table 3,Avellaneda et al. (2022))
@@ 6.1 IS–LM foundations for (multi-)torus dynamics: a mathematical sketch
1. IS–LM(+Phillips) linear core and oscillatory modes,Hicks (1937)
@@
The $(y,r)$-subsystem has complex eigenvalues (oscillatory adjustment,Benhabib and Nishimura (1979); Barnett and Duzhak (2008)) iff
@@
Note on “Hopf”.Kuznetsov (2013)
@@ 3. From one torus to several: a multi-block macro
Writing $z^{(s)}=\rho_{s}e^{i\theta^{(s)}}$ and projecting on the limit cycles ($\rho_{s}\approx\rho_{s}^{\star}$), we obtain a phase network,Ikeda et al. (2012),
@@ 6.2 Future research agenda (testable directions)
Combine GIM with mixed-frequency nowcasting: update $(\theta_{t},\phi_{t})$ and the tangent-space forecasts as high-frequency data arrive, and reweight sleeves by expanding-SVD eigenvalues and curvature gates. This evaluates whether real-time macro signals enhance geometry-aware timing.Gutiérrez-Fandiño et al. (2021)
@@ Appendix C Scenario catalog
Appendix D Coordinate transforms,do Carmo (2016, 1992)